OpenStudy (anonymous):

(4+2x)(7+2x) = 238 find x and show all work please:)

OpenStudy (anonymous):

x=5

OpenStudy (anonymous):

how do you find that answer?

OpenStudy (anonymous):

28 +22x+4x2 =238 4x2+22x-210 = 0 divide with 2 2x2+11x-105=0 (2x+21)(x-5) =0 x = -21/2 , x=5

OpenStudy (anonymous):

WolframAlpha.com